OpenStudy (anonymous):

Jason can run 7 3/7 miles in one hour. how long will it take Jason to run nine laps on the school's quarter-mile track?

OpenStudy (anonymous):

hi :P

OpenStudy (anonymous):

i know the answer

OpenStudy (nowhereman):

If you assume he runs that distance at the same speed you only divide 9/4 through 7 3/7 to get the time in hours.

OpenStudy (anonymous):

the answer is in my tight buthole

OpenStudy (nowhereman):

But I would guess, he could run faster, as he does not have to run as far ;-)

OpenStudy (anonymous):

wait whatttt! hahah im confoooosedd! is the answer .3?

OpenStudy (nowhereman):

Approximately it is .3 yes, exactly it is \[\frac{9}{4} : 7\frac 3 7 = \frac{63}{208}\]

OpenStudy (anonymous):

oh ok! thank you :-)